1. Concentrated sulphuric acid is a solution that has a density of 1.84 g/ml and contains 98.3


% H2SO4 by mass. What is the molarity of this acid?


2. What is the molality of a solution that contains 0.850 g of ammonia, in 125 g of water?


3. How would you distinguish molarity, molality and normality?

1
Expert's answer
2021-11-02T02:23:56-0400

Molar mass of NH3= 17.031

0.85/17.031= 0.0499mol

= 0.0499/0.125=0.399mol/L

Molarity ( ) is defined as the number of moles of solute per liter of solution. Molality ( ) is defined as the number of moles of solute per kilogram of solvent. Normality ( ) is defined as the number of equivalents per liter of solution.